Activity at the Soufriere Hills volcano has been at a relatively low level throughout the past week, with only one area of slow dome growth noted in the northern part of the active area within English's Crater. Seismicity has been at a very low level and there has been no measurable deformation of the volcanic edifice.
Seismic activity currently comprises four main elements. Volcano-tectonic earthquakes continue at a rate of 2 to 3 per day, with an exception being 1 January when 17 were recorded in an eight hour period following a regional earthquake north of Trinidad (see below). These events have locations somewhat scattered around the Soufriere Hills volcano at depths of 0 to 7 km; the 1 January swarm clustered just north of the active crater area at depths of 1 to 3 km. Long-period events recorded at most seismic stations have been occurring at a rate of 10 to 15 per day; these cannot be located but are thought to source from very shallow depths in the crater area. Broadband tremor has been recorded intermittently on the Gages seismic station throughout the week; it is of consistently low amplitude and is thought to be due to variable levels of steam venting from the crater area. This tremor was essentially continuous for most of 28 December.
A total of 11 regional earthquakes have been recorded by the local network during this week. The most notable of these was a magnitude 5.0 event located at a depth of 25 km approximately 55 km north of Port of Spain, Trinidad. This earthquake was felt as far north as Martinique, and may have been the trigger for the small swarm of volcano-tectonic earthquakes beneath the Soufriere Hills volcano.
EDM lines around the volcano continue to be measured on a routine basis. None of the lines showed any changes during the week beyond the level of accuracy of the measurements. GPS sites were removed on 29 December due to commitment of the equipment for other projects; permanent sites are expected to be installed during the second half of January. The data collected during the last 10 days of the GPS survey have not yet been processed, but the results are expected by next week. The Long Ground tiltmeter continues to show no measurable deformation.
Visual observations have been possible throughout the week from a number of vantage points. Dome growth on the eastern side of the active area slowed to a stop during the week; 3 m of vertical growth was recorded over the period 28 to 30 December accompanied by rock avalanching from the flanks, but both growth and rock avalanching had essentially stopped by 1 January. A second active area at the northern edge of Castle Peak Dome, which was first noted by the presence of a small spine on 26 December, has grown in volume throughout the week, although its vertical height has remained constant. This dome is only visible on the ground from the Harris' area, and visibility into the area from the air has been restricted by steam generation close by. It appears, however, that rock spalling is occurring into the moat area from this area of active growth, although no material has moved further down the moat to the east.
Only one notable ash-generating event occurred during this period; a plume of ash drifted westward late in the afternoon on 29 December following a substantial rock avalanche in the crater area. A continuous steam plume has been present, sometimes charged with small amounts of ash. Steam is being emitted from a number of areas within the crater.
An attempt to collect rock samples from the new dome was unsuccessful; further attempts are planned for the coming week.